OpenStudy (anonymous):

How would you work this out? 2/3x = -2/3x +4 I haven't worked with fractions (this way) before, so I'm not sure how to do it.

pooja195 (pooja195):

Are we isolating x?

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

Ok we we want x's on one side and numbers without the x on another Start by adding 2/3's to both sides 2/3x+2/3x=-2/3x+2/3+4

OpenStudy (michele_laino):

hint: if I add \((2/3)x\) to both sides, I can write: \[\frac{2}{3}x + \frac{2}{3}x = - \frac{2}{3}x + 4 + \frac{2}{3}x\]

pooja195 (pooja195):